OpenStudy (anonymous):

how do you solve 2cos(z)cot(z) = 3 ??? If someone can just show the working till it becomes either sin or cos, thats cool :P

OpenStudy (anonymous):

cot= cos/sin cos*cot=cos^2/sin= (1-sin^2)/sin now, make the quadratic..

OpenStudy (callisto):

Give you something more pretty :) 2cos(z)cot(z) = 3 2cos(z)[cos(z)/sin(z)] = 3 2[cos(z)]^2 = 3sin(z) 2[1-(sinz)^2] = 3sinz 2-2(sinz)^2 - 3sinz =0 2(sinz)^2 + 3sinz -2 =0 Solve the quadratic and work out the answer :)

OpenStudy (anonymous):

ohhhhhhh. ha, ha, ha, i feel so stupid right now, its not funny -_-